OpenStudy (anonymous):

An employee's salary is calculated by the formula S(t)=S0 e^.08t. t=0 corresponds to the beginning of 2011. in 2014 this employee was paid $69,918.70. What was the employee's annual percent salary increase?

OpenStudy (tkhunny):

How many years have passed?

OpenStudy (anonymous):

t=0 at the beginning of 2011. The salary $69,918.70 is in the year 2014, so 3 years.

OpenStudy (tkhunny):

All right. What, then is \(A_{0}\)?

OpenStudy (anonymous):

That is what I'm trying to solve, isn't it? When I went through what I thought was the process, I got an answer of 46. something, which is definitely not the answer.

OpenStudy (tkhunny):

Of course, I meant \(S_{0}\). You have: \(S(t)=S_{0}e^{.08t}\) You have: \(t = 3\) You have: \(S(3) = 69,918.70\) Solve for \(S_{0}\)

OpenStudy (tkhunny):

\(69918.70 = S_{0}e^{0.08(3)}\) -- Solve for \(S_{0}\).

OpenStudy (anonymous):

I must be doing something wrong. Here's my process: \[S(t)=S _{0}e ^{0.08t}\] \[69918.7=S _{0}e ^{0.08(3)}\] \[69918.7=S _{0}e ^{.24}\] \[\ln 69918.7=S _{0}(.24)\ln e\] with the natural log canceling out e on the right side \[11.155=S _{0}(.24)\] Divide both sides by .24 and I get \[S _{0}=46.4795\]

OpenStudy (anonymous):

however, if I don't use natural log and just divide by e^.24, I get S0= 54999.99. Is this correct? If so, why would I not use natural log?

OpenStudy (tkhunny):

Try: \(\ln(69918.70) = \ln(S_{0}) + 0.24\) How did \(S_{0}\) manage to avoid the logarithm?